# Perform Professional CRM Tasks in Realistic Environments

### v1

@@ Abstract
Customer Relationship Management (CRM) systems are vital for modern enterprises, providing a foundation for managing customer interactions and data. Integrating AI agents into CRM systems can automate routine processes and enhance personalized service. However, deploying and evaluating these agents is challenging due to the lack of realistic benchmarks that reflect the complexity of real-world CRM tasks. To address this issue, we introduce CRMArena, a novel benchmark designed to evaluate AI agents on realistic tasks grounded on professional work environments. We worked with CRM experts to design nine customer service tasks distributed across three personas: service agent, analyst, and manager. We synthesize a large-scale simulated organization, populating 16 commonly-used industrial objects (e.g., account, order, knowledge article, case) with high interconnectivity, and uploading it into a real Salesforce CRM organization. UI and API access to the CRM is provided to systems that attempt to complete the tasks in CRMArena. Experimental results reveal that state-of-the-art LLM agents succeed in less than 40% of the tasks with ReAct prompting, and less than 55% even when provided manually-crafted function-calling tools. Our findings highlight the need for enhanced agent capabilities in function-calling and rule-following to be deployed in real-world work environment. CRMArena is an open challenge to the community: systems that can reliably complete tasks showcase direct business value in a popular work environment. [^1] Our code and benchmark have been released at https://github.com/SalesforceAIResearch/CRMArena.
@@ 3 Benchmarking Experiments
3.1 Experimental Settings
Models
We evaluate state-of-the-art proprietary and open-source LLMs, including gpt models (gpt-4o and gpt-3.5-turbo); claude models (claude-3.5-sonnet and claude-3-sonnet), and the llama models (llama-3.1-405b and llama-3.1-70B).  [^8] We observed the native function-calling mode to perform poorly and hence report the prompt mode performance for Llama 3.1 models. With these models, we tested three common agentic frameworks: Act, ReAct Yao et al. (2023), and Function Calling (FC). ReAct is a prompt-based method, with each step characterized by a thought and action process, while Act is ReAct without the thought component. The details of these settings are described in the following paragraphs and Appendix G.
Action Space
[…]
Observation Space
Evaluation Metrics
@@ 3.2 Results
The main results are summarized in Table 2. We made the following observations. First, real-world CRM tasks remain challenging for top LLM agents. Using the ReAct framework, the best model (gpt-4o) only achieves an overall score of 38.2%. Even when equipped with human-crafted functions, the overall performance is still only 54.4%. These findings highlight the challenges of our CRMArena. Second, stronger and weaker LLMs show opposite trend on different agentic frameworks. In particular, models like gpt-4o and claude-3.5-sonnet score higher in the FC setting, while their weaker counterparts performs worse when equipped with function calling capabilities. This indicate that human-defined functions may not always help LLM agents, as weaker models may not be able to properly utilize the functions, resulting in lower performance. Finally, open-source models are catching up the proprietary LLMs. Across three settings, we see the llama models score similar, and sometimes higher, than the gpt and claude models. This indicate a closing gap between the open and closed-source models. From Figure 6, we observe how llama models tend to show higher scope for error recovery based on execution feedback than the closed-source models.
Model # Completion Tokens # Turns Cost (\$) ReAct gpt-4o 48568.73 5.4 0.182 claude-3.5-sonnet 70814.75 6.9 0.228 llama3.1-405b 35647.29 7.3 0.125 FC gpt-4o 78305.38 6.8 0.305 claude-3.5-sonnet 105248.43 8.1 0.371
@@ 3.3 Discussions
In two-thrid of the agentic frameworks, gpt-4o performs the best. The efficiency of gpt-4o is also reflected in Table 3, which shows that gpt-4o has the lowest cost per instance and requires the least number of turns to complete a query. Therefore, the most cost-effective solution is using gpt-4o under the function calling setting.
@@ 4 Related Work
Several benchmarks have been developed to evaluate LLM-based agents Yao et al. (2022); Liu et al. (2024); Jimenez et al. (2024). Recently, major efforts have focused specifically on web agents, which challenges LLMs to navigate and perform actions on websites. These websites are often about everyday scenarios, such as e-commerce, and social discussion form Deng et al. (2023); He et al. (2024); Zhou et al. (2024); Lù et al. (2024); Yoran et al. (2024). Another line of work focus on evaluating the safety of deploying agents Ruan et al. (2024); Yuan et al. (2024).

### v2

@@ Abstract
Customer Relationship Management (CRM) systems are vital for modern enterprises, providing a foundation for managing customer interactions and data. Integrating AI agents into CRM systems can automate routine processes and enhance personalized service. However, deploying and evaluating these agents is challenging due to the lack of realistic benchmarks that reflect the complexity of real-world CRM tasks. To address this issue, we introduce CRMArena, a novel benchmark designed to evaluate AI agents on realistic tasks grounded on professional work environments. We worked with CRM experts to design nine customer service tasks distributed across three personas: service agent, analyst, and manager. We synthesize a large-scale simulated organization, populating 16 commonly-used industrial objects (e.g., account, order, knowledge article, case) with high interconnectivity, and uploading it into a real Salesforce CRM organization. UI and API access to the CRM is provided to systems that attempt to complete the tasks in CRMArena. Experimental results reveal that state-of-the-art LLM agents succeed in less than 58% of the tasks with ReAct prompting, and less than 65% even when provided manually-crafted function-calling tools. Our findings highlight the need for enhanced agent capabilities in function-calling and rule-following to be deployed in real-world work environment. CRMArena is an open challenge to the community: systems that can reliably complete tasks showcase direct business value in a popular work environment. [^1] Our code and benchmark have been released at https://github.com/SalesforceAIResearch/CRMArena.
@@ 3 Benchmarking Experiments
3.1 Experimental Settings
Models
We evaluate state-of-the-art proprietary and open-source LLMs, including gpt models (gpt-4o and gpt-3.5-turbo); claude models (claude-3.5-sonnet and claude-3-sonnet), and the llama models (llama-3.1-405b and llama-3.1-70B Dubey et al. (2024)). [^8] We observed the native function-calling mode to perform poorly and hence report the prompt mode performance for Llama 3.1 models. Additionally, we tested inference-time scaling models for enhanced reasoning capabilities, including o1 and deepseek-r1 Guo et al. (2025). With these models, we tested three common agentic frameworks: Act, ReAct Yao et al. (2023), and Function Calling (FC). ReAct is a prompt-based method, with each step characterized by a thought and action process, while Act is ReAct without the thought component. The details of these settings are described in the following paragraphs and Appendix G.
Action Space
[…]
Observation Space
Evaluation Metrics
@@ 3.2 Results
The main results are summarized in Table 2. We made the following observations. First, real-world CRM tasks remain challenging for top LLM agents. Using the ReAct framework, the best model (o1) only achieves an overall score of 57.7%. Even when equipped with human-crafted functions, the overall performance is still only 64.3%. These findings highlight the challenges of our CRMArena. Second, stronger and weaker LLMs show opposite trend on different agentic frameworks. In particular, models like gpt-4o and claude-3.5-sonnet score higher in the FC setting, while their weaker counterparts performs worse when equipped with function calling capabilities. This indicate that human-defined functions may not always help LLM agents, as weaker models may not be able to properly utilize the functions, resulting in lower performance. An intriguing exception is deepseek-r1. Though deepseek-r1 is recognized as a strong reasoning model, its tool-calling capabilities seem lacking, primarily due to its (1) inadequate adherence to user instructions and (2) poor ability to adjust previous responses based on external feedback. function calling might be unnecessary with a sufficiently strong reasoning model, as evidenced by o1 in the ReAct setting outperforming all other models in the FC setting. Nevertheless, integrating human-crafted functions can still offer performance benefits to strong reasoning models like o1. Finally, open-source models are catching up the proprietary LLMs. Across three settings, we see the llama models score similar, and sometimes higher, than the gpt and claude models. This indicate a closing gap between the open and closed-source models. From Figure 6, we observe how llama models tend to show higher scope for error recovery based on execution feedback than the closed-source models.
Model # Completion Tokens # Turns Cost (\$) ReAct gpt-4o 48,568.73 5.4 0.182 claude-3.5-sonnet 70,814.75 6.9 0.228 llama3.1-405b 35,647.29 7.3 0.125 FC gpt-4o 78,305.38 6.8 0.305 claude-3.5-sonnet 105,248.43 8.1 0.371
@@ 3.3 Discussions
Excluding the two reasoning models, in two-third of the agentic frameworks, gpt-4o performs the best. The efficiency of gpt-4o is also reflected in Table 3, which shows that gpt-4o has the lowest cost per instance and requires the least number of turns to complete a query. Therefore, the most cost-effective solution is using gpt-4o under the function calling setting.
@@ 4 Related Work
Several benchmarks have been developed to evaluate LLM-based agents Yao et al. (2022); Liu et al. (2024); Jimenez et al. (2024). Recently, major efforts have focused specifically on web agents, which challenges LLMs to navigate and perform actions on websites. These websites are often about everyday scenarios, such as e-commerce, and social discussion form Deng et al. (2023); He et al. (2024); Zhou et al. (2024); Lù et al. (2024); Yoran et al. (2024). Another line of work focus on evaluating the safety of deploying agents Ruan et al. (2024); Yuan et al. (2024); Yin et al. (2024); Qiu et al. (2025).
@@ Appendix G Implementation Details
  • llama3.1-8b: meta-llama/Meta-Llama-3.1-8B-Instruct-Turbo